# Recurrent Diffusion for Large-Scale Parameter Generation

### v1

@@ Abstract
Parameter generation has struggled to scale up for a long time, significantly limiting its range of applications. In this study, we introduce Recurrent diffusion for large-scale Parameter Generation, called RPG. We first divide the trained parameters into non-overlapping parts, after which a recurrent model is proposed to learn their relationships. The recurrent model’s outputs, as conditions, are then fed into a diffusion model to generate the neural network parameters. Using only a single GPU, recurrent diffusion enables us to generate popular vision and language models such as ConvNeXt-L and LoRA parameters of LLaMA-7B. Meanwhile, across various architectures and tasks, the generated parameters consistently perform comparable results over trained networks. Notably, our approach also shows the potential to generate models for handling unseen tasks, which largely increases the practicality of parameter generation. Our code is available here.
Kai and Dongwen contributed equally.
@@ 1 Introduction
Looking back on the journey of deep learning, the scaling up of neural networks is one of the most important keys to its remarkable success across various tasks [33, 22, 44]. In contrast, neural network parameter generation, from HyperNetworks [21] to recent diffusion-based methods [51, 69, 66], has struggled to scale up effectively, limiting its practical applications. As shown in Fig. 1, the scale gap between vision (or language) models and the generated parameters is at least $\mathbf{10^{4}}$, posing significant challenges for this field.
Figure 1: Partial roadmap of vision, language, and parameter generation models. Parameter number in vision or language models is at least $\mathbf{10^{4}}$ times larger than that of generated parameters.
To figure out the key challenges in scaling up parameter generation, we first analyze its unique requirements. Unlike traditional deep learning models that typically process data such as images or text, parameter generation involves network parameters in the training process. The size of parameters could be significantly larger than images or texts size. This fundamental difference in input format presents a significant challenge when scaling up. As the size of generated parameters increases, the GPU memory requirement quickly becomes prohibitive.
Recently, p-diff [69] and D2NWG [66] attempted to address the challenge of balancing memory constraints with large-scale parameter generation. P-diff mainly generates a subset of the entire neural network parameters, while D2NWG employs a two-step process of synthesizing parameter parts and then combining them to form a complete model. However, these methods may overlook the inherent correlations among the parameter parts. To study the impact of these correlations, we conduct an experiment: exchanging partial parameters between two ViT-Tiny [14] models trained on CIFAR-10 [32] dataset. As illustrated in the following table, the significant performance degradation highlights the critical importance of parameter correlations.
@@
How can we model parameter relationships and leverage them for efficient parameter generation? Vision transformers often outperform CNNs by modeling patch relationships and capturing global information via self-attention. In language tasks, LLMs use next token prediction to model token relationships, capturing long-range dependencies. Inspired by these approaches, we can consider treating parameter parts as tokens in neural network parameter generation, potentially enabling methods to model inter-parameter relationships and capture dependencies.
In this paper, we propose Recurrent diffusion for large-scale neural network Parameters Generation (RPG). Our approach first divides the trained network parameters into a set of non-overlapping parameter parts (simply called ‘tokens’ in the following). Subsequently, we use a recurrent model to learn the relationships among the tokens. Finally, the outputs of the recurrent model, as conditions, are fed into a diffusion process to generate the network parameters.
Our approach has the following properties: i) With a single GPU, our approach successfully synthesizes full parameters of large-scale vision models such as ResNet [22], ViT [14], and ConvNeXt [41] series, as well as LoRA [27] parameters of LLaMA-7B [67]. ii) Across various architectures and tasks, the generated parameters achieve comparable performance to the original models while exhibiting high diversity. iii) Empirical evidence suggests our approach has potential for generating models in unseen tasks. We anticipate that this work will inspire future research in large-scale parameter generation.
2 Related Works
Diffusion models.
Diffusion models [23, 48, 13] gain increasing popularity in recent years, due to their superiority in image generation. Ever since its advent, many works have been done focusing on improving the generation quality and efficiency of diffusion models. For generation quality, Rombach et al. [55] propose to conduct diffusion in the latent space, enabling high-resolution image synthesis. Peebles and Xie [50] leverage the transformer [68] to explore scalability of diffusion models, proving the possibility of generating higher quality images by increasing model size. As for efficiency problem, efficient samplers [64, 42, 65], efficiency models [16, 62, 74], and global acceleration approaches [43, 49] are proposed to increase diffusion models’ efficiency. These methods facilitate generating high quality images with less computational and/or memory cost. Although diffusion models for image generation have achieved great success, how to improve quality and efficiency in large-scale parameter generation remains to be explored.
Recurrent models.
Recurrent neural networks (RNNs) were first proposed to process sequential data, i.e., the text. To tackle the vanishing gradient problem in early RNNs, long short-term memory (LSTM) [24, 25]. In recent years, transformer-based models [68] exhibits excellent potential in sequential data processing, due to their parallelized training and scalability. Although most transformers [53, 67] are used in auto-regressive manner, they can be seamlessly adapted to the recurrent paradigm. However, transformer-based models are suffering from the the quadratic complexity during their generation process, which greatly hinders their efficiency. Recently, various attempts, such as linear attentions [70, 7], RWKV [52], Mamba [19, 11], and xLSTM [2], have been proposed to tackle this problem.
Parameter generation.
The core idea of parameter generation is to learn the distribution of trained parameters. Stochastic neural networks [63, 5, 71, 57, 46, 18] and Bayesian neural networks [47, 30, 54, 31, 4, 17] model the priors or probability distributions over the parameters. These approaches mainly employ the learned prior knowledge of parameters to improve robustness and generalization, and to mitigate overfitting and uncertainties. However, these methods are limited by their generality to large-scale generation or more complex real-world scenarios.
HyperNetworks [21], i.e., a small network, is proposed to generate various architectures’ parameters for a larger network. Smash [6] extends the range of architectures via a memory read-writes scheme. With the development of diffusion models, many works [51, 8, 15, 69, 66, 38, 37, 29] adopt diffusion models to generate neural network parameters. G.pt [51] collects 23 million checkpoints as the training data and uses conditional diffusion to learn the distribution of the parameters. Besides the heavy cost of collecting model parameters, G.pt can only generate less than 10K parameters. P-diff proposes unconditional diffusion to simulate the parameter updating. HyperRep-resentations [58, 59, 61] use an autoencoder to capture the latent distribution of trained models. COND P-DIFF [29] and Tina [37] introduce task-controlled or text-controlled parameter generation method. Unfortunately, the above methods have a common drawback: can not generate large-scale parameters, such as whole parameters of ResNet, ViT, ConvNeXt, or LoRA. Therefore, our approach brings new inspiration to the field of parameter generation.
3 Large-Scale Parameter Generation
@@ 3.1 Overview
Our approach comprises two key components: parameter tokenization and recurrent diffusion. We show the inference process of recurrent diffusion in Fig. 2. The permutation state and position embedding are fed into the recurrent model. Then, the outputs of the recurrent model serve as conditions for the diffusion process, which generates the entire neural network parameters.
Figure 2: Illustration of the inference of the recurrent diffusion. The recurrent model takes permutation state and position embedding as inputs. The outputs of the recurrent model are then fed into the diffusion process as condition to synthesize the neural network parameters.
3.2 Parameter Tokenization
Inspired by the success of language and vision models [68, 14], we propose parameter tokenization that divides network parameters into non-overlapping tokens. Considering the distribution shifts across different layers, we first categorize the trained parameters according to their respective layer indices. Then, we apply normalization (subtracting the mean and dividing by the standard deviation) on each layer. These operations can be formulated as follows,
@@ 3.2 Parameter Tokenization
where $W$ denotes the trained parameters. $\mu$ and $\sigma$ denote the mean and standard deviation values of parameters. $w_{i}$ and $\hat{w}_{i}$ are the original and normalized parameters of the $i$-th layer, respectively.
The number of parameters varies across these layers, which is not conducive to efficient batch training. To this end, we slice each layer parameter into a set of tokens with the same token size, which can be written as follows,
@@
where the $k_{i}^{j}$ represents the $j$-th token of $i$-th layer. For the last token of each layer, we apply padding operation to ensure that all layers have tokens of uniform length. It is worth noting that the padded regions are excluded from the loss calculation.
@@ 3.3 Recurrent Diffusion
Permutation state.
Neural network symmetries [1, 34] do not affect the model outcomes but increase the difficulty of learning the parameter distribution. To address this, we introduce a unique state for each trained model $W$ via one-hot embedding. This operation provides a guide for the generated parameters to mitigate the influence of parameter symmetries. For simplicity, we use $S$ to represent the permutation state of $W$.
Position embedding.
Inspired by ViT [14], we also encode the layer and token information described in the parameter tokenization (Sec. 3.2) using a two-dimensional sinusoidal position encoding. Specifically, the first dimension encodes the layer index of the token in the original model, while the second dimension encodes the position of the token within its layer. For $i$-th layer parameter tokens $K[i]$, the position embedding can be formulated as follows,
[…]
where $e_{i}^{j}$ denotes the position embedding of the parameters belong to $j$-th token of $i$-th layer.
@@
After obtaining the parameter tokens, permutation states, and position embeddings, we use a recurrent model to learn the representation of the parameter tokens. For clarity, we will refer to the output of the recurrent model as the ‘prototype’ in the following. This operation can be written as follows:
$P_{i}^{j},H_{i}^{j}=f(H_{i}^{j-1},e_{i}^{j},S),i\in[1,I],j\in[1,J],$ (4)
where $P_{i}^{j}$ and $H_{i}^{j}$ denote the prototype and hidden state of the parameters belonging to $j$-th token of $i$-th layer, respectively. $f(\cdot,\cdot,\cdot)$ denotes the state transition function. The structure of the recurrent model is simple. Considering efficiency, we use Mamba [20] followed by an MLP to project the feature dimension to the required size for the diffusion model. We also conduct ablation studies with other recurrent model architectures, such as LSTM [26] and transformer with its decoder in a causal manner [68].
@@
Inspired by p-diff [69] and MAR [35], we use 1D convolution to build the diffusion model. In this part, the parameter prototypes, serving as conditions, are fed into the diffusion process along with random noise. We optimize our approach through the following equation:
$L_{\text{diff}}=\mathbb{E}_{t,K,\epsilon}[||\epsilon-\epsilon_{\theta}(K_{t},t,P)||^{2}],$ (5)
where $K$, $P$, $L_{\text{diff}}$, and $t$ denote parameter tokens, prototypes, diffusion loss, and time step, respectively. $\epsilon$ is Gaussian noise and $\epsilon_{\theta}$ is the denoising network parameterized by $\theta$. Note that, the gradient propagates through $P$ to the recurrent model, implicitly optimizing it as well.
4 Experiments
@@ 4.1 Setup
We mainly evaluate our method across a wide range of tasks, including ImageNet-1K [12] for the classification, ADE20K [77] for the semantic segmentation, COCO [39] for the object detection, and BoolQ [9], PIQA [3], SIQA [56], HellaSwag [75], and ARC [10] for the commonsense reasoning tasks. To verify the scalability of our approach, we conduct experiments on various architectures with parameter counts ranging from several to hundred million. Details of parameter counts can be found in Tab. 1, 2, 3.
@@
We take parameters collection on the ImageNet-1K as an example. To save the cost, we finetune the full parameters of the models released in timm  [^1] https://github.com/huggingface/pytorch-image-models and save 50 checkpoints as the training data. For each checkpoint, we assign a unique permutation state to guide the generated parameters.
@@
Training details.
We default to using Mamba [20] as the architecture of the recurrent model. The length of parameter tokens, permutation states, position embeddings, and prototypes is set to 8192. It is worth noting that the permutation states and position embeddings are fixed during the training by default. We also study the influence of the token length, varying it from 1024 to 16384. The parameter diffusion consists of 1D convolutional layers. More details about the model architectures, hyperparameters, and training process can be found in Appendix B.1.
@@
We input permutation states and position embeddings into the recurrent model to generate the prototypes. Then, the diffusion model utilizes the prototypes as conditions, along with random noises, to synthesize the entire network parameters. We repeat the above process 10 times and report the best, average, and minimum.

### v2

# Scaling Up Parameter Generation: A Recurrent Diffusion Approach

@@ Abstract
Parameter generation has long struggled to match the scale of today’s large vision and language models, curbing its broader utility. In this paper, we introduce Recurrent Diffusion for Large-Scale Parameter Generation (RPG), a novel framework that generates full neural network parameters—up to hundreds of millions—on a single GPU. Our approach first partitions a network’s parameters into non-overlapping ‘tokens’, each corresponding to a distinct portion of the model. A recurrent mechanism then learns the inter-token relationships, producing ‘prototypes’ which serve as conditions for a diffusion process that ultimately synthesizes the full parameters. Across a spectrum of architectures and tasks—including ResNets, ConvNeXts and ViTs on ImageNet-1K and COCO, and even LoRA-based LLMs—RPG achieves performance on par with fully trained networks while avoiding excessive memory overhead. Notably, it generalizes beyond its training set to generate valid parameters for previously unseen tasks, highlighting its flexibility in dynamic and open-ended scenarios. By overcoming the longstanding memory and scalability barriers, RPG serves as a critical advance in ‘AI generating AI’, potentially enabling efficient weight generation at scales previously deemed infeasible.
Machine Learning, ICML
1National University of Singapore  2University of St.Gallen  3University of Texas at Austin
00footnotetext: equal contribution, corresponding author11footnotetext:  Code: NUS-HPC-AI-Lab/RPG
@@ 1 Introduction
Scaling up neural networks has been one of the most crucial factors driving the progress of deep learning, enabling remarkable performance across a wide range of tasks (Krizhevsky et al., 2012; He et al., 2016; Meta, 2024). However, neural network parameter generation—from early HyperNetworks (Ha et al., 2017) to more recent diffusion-based methods (Peebles et al., 2022; Wang et al., 2024; Soro et al., 2024)—has not kept pace with the expansion of mainstream vision and language models. Indeed, as illustrated in Fig. 1, there is an astonishing scale gap of at least $10^{3}$ between the parameter counts of widely used models (e.g., ViT, ConvNeXt, LLaMA) and the largest parameter generators available today. This discrepancy not only underscores the challenge of scalability but also confines existing generators to academic demonstrations rather than broad, real-world applicability.
Figure 1: Partial roadmap of vision, language, and parameter generation models. Parameter number in vision or language models is at least $\mathbf{10^{3}}$ times larger than that of generated parameters.
Traditional approaches such as HyperNetworks struggle to handle even moderately sized models, primarily due to memory overheads and optimization complexity. While recent diffusion-based works attempt to break the memory bottleneck via partial-parameter or two-stage generation (Wang et al., 2024; Soro et al., 2024), such methods may overlook key correlations among different parameter subsets. To illustrate the importance of capturing inter-part correlations, we conduct a simple experiment: we train two ViT-Tiny (Dosovitskiy et al., 2021) models (A and B) on CIFAR-10 (Krizhevsky & Hinton, 2009), then exchange half of their layer parameters. Despite both models performing well originally (i.e., $90.4\%$ for model A and $89.6\%$ for model B), the hybrid model’s accuracy drops dramatically to $45.8\%$ (as seen below).
@@
This striking performance gap highlights that adequately modeling parameter correlations is not a trivial detail: a robust generator must therefore explicitly model such relationships while also surmounting the memory pitfalls of large-scale processing.
Our contributions.
We present Recurrent Diffusion for Large-Scale Parameter Generation (RPG), a novel, end-to-end framework capable of generating full neural network parameters with up to hundreds of millions of weights using just a single commodity GPU. To the best of our knowledge, RPG is the first method to efficiently synthesize the parameters of models like ConvNeXt-L and LLaMA-LoRA–with up to $\sim$200M weights–in a single pass. This significant leap is achieved by two core technical innovations:
  1. 1.
    Parameter processing for large-scale tokenization. We devise a highly customized parameter tokenization strategy designed to preserve both the layer-wise distribution and the cross-layer correlations. Specifically, we (i) split weights according to their layer indices, (ii) apply layer-wise normalization to mitigate distribution shifts, (iii) slice parameters into non-overlapping tokens with uniform size, and (iv) introduce a lightweight permutation state to alleviate symmetry issues when collecting multiple checkpoints. Additionally, we employ 2D position embeddings (layer index + token index) to ensure the network retains positional awareness of each token within the entire set. This tailored pre-processing pipeline is crucial for learning stable and high-quality parameter representations.
  2. 2.
    Recurrent diffusion modeling. Unlike prior diffusion-based generators that flatten or divide all parameters into a single sequence or chunks, we propose a recurrent mechanism to learn the relationships among tokens. Concretely, we map each token into a hidden space via a recurrent model, whose outputs serve as prototypes. These prototypes then condition a 1D diffusion process that progressively denoises the parameter tokens. By recurrently modeling global inter-token interactions and leveraging a diffusion sampler for synthesis, RPG aligns well with the dual goals of capturing token-to-token dependencies and remaining memory-efficient.
Key results and impact.
Extensive evaluation on Image-Net classification, ADE20K segmentation, COCO detection, and commonsense reasoning show that RPG consistently produces weights on par with trained models. Notably:
  • Single-GPU feasibility. Our recurrent design and parameter-token strategy enable inference on a single commodity GPU beyond 100M weights.
  • Generalization to unseen tasks. RPG can generate high-performing networks for unseen binary classification tasks on CIFAR-10, providing evidence for its broader generalization capabilities.
  • Architectural Versatility. Our method supports diverse families such as ResNet, ViT, ConvNeXt, and LoRA-based language models, making it a unified framework for parameter generation.
These results place us significantly closer to the vision of ‘AI creating AI’, where a single generative model can synthesize entire networks tailored for different tasks. We believe RPG’s strong scalability, lightweight memory footprint, and ability to accurately capture global parameter relationships may open new frontiers in fields like rapid architecture search, efficient model adaptation, and beyond.
2 Method: Large-Scale Parameter Generation
Figure 2: Illustration of parameter processing (left) and inference of recurrent diffusion (right). The recurrent model integrates permutation states and position embeddings, generating prototypes that condition the diffusion model to synthesize the full parameters.
In this section, we describe our proposed Recurrent Diffusion for Large-Scale Parameter Generation (RPG), which synthesizes full neural network parameters under strict memory constraints. As illustrated in Fig. 2, our approach comprises two primary components: parameter processing and recurrent diffusion.
Parameter processing (Sec. 2.2) equips the model with a specialized tokenization pipeline, layer-wise normalization, permutation states, and position embeddings—all meticulously designed to mitigate issues like distribution shifts, neural symmetry, and the sheer size of modern network weights. Recurrent diffusion (Sec. 2.3) then learns these ‘tokens’ via a recurrent model whose outputs condition a 1D diffusion process, enabling us to capture global parameter correlations while maintaining feasibility on a single GPU.
@@ 2.1 Overview
Our pipeline for large-scale parameter generation is shown conceptually on the right of Fig. 2. Given a set of trained parameters $\{W\}$ (e.g., from various checkpoints or model variants), we first transform them into parameter tokens, augment them with a permutation state, and apply position embeddings. This processed sequence is fed into a recurrent model, which outputs per-token ‘prototypes’. Finally, a diffusion model uses these prototypes, along with random noise, to synthesize the entire parameter vector. In practice, we can sample from this generative model to produce a diverse family of high-performing weights.
2.2 Parameter Processing
Parameter tokenization.
Drawing inspiration from the success of language and vision transformers (Vaswani, 2017; Dosovitskiy et al., 2021), we divide a network’s parameters into a sequence of non-overlapping tokens. However, direct flattening or naive partitioning often fails to preserve crucial layer-wise statistics and undermines global consistency. To circumvent this, we first segregate parameters by layer and then normalize each layer to reduce distribution shifts:
@@ 2.2 Parameter Processing
where $w[i]$ and $\hat{w}[i]$ denote the original and normalized parameter blocks of the $i$-th layer, respectively, and $\mu,\sigma$ are the mean and standard deviation of that layer’s weights.
Next, each normalized layer $\hat{w}[i]$ is tokenized into contiguous chunks of uniform size $k$ (plus padding if necessary):
@@
Importantly, the padded regions are excluded from the loss and do not affect gradient updates. By enforcing a consistent token size across layers, we facilitate batch processing of parameter tokens—a critical step for large-scale training.
Permutation state.
When aggregating checkpoints from various training runs, neural symmetry (Kunin et al., 2021) can become a major bottleneck. Essentially, distinct permutations of layer weights may yield identical final performance but appear different to a generative model. To resolve this ambiguity, we introduce a distinct permutation state $S$—encoded as a one-hot vector—for each checkpoint $W$. This extra condition guides the generative model to treat otherwise symmetrical weight configurations as unique exemplars, significantly stabilizing distribution learning.
Position embedding.
To preserve token locality and inter-layer relationships, we augment each token with a 2D sinusoidal position embedding (Dosovitskiy et al., 2021). Specifically, for the $j$-th token of layer $i$, we assign
@@
The first dimension encodes the layer index, while the second dimension captures the token’s in-layer position. As a result, the g model knows not only how tokens are grouped by layer, but also each token’s placement within that layer.
@@ 2.3 Recurrent Diffusion
After obtaining the parameter tokens $K[i]$, permutation states $S$, and position embeddings $e[i]$, we feed them into a recurrent network $f(\cdot)$ that learns token-wise representations while capturing cross-token dependencies. We denote the output as a prototype $P_{i}^{j}$:
$P_{i}^{j},\,H_{i}^{j}\;=\;f\!\Bigl{(}H_{i}^{j-1},\,e_{i}^{j},\,S\Bigr{)},$ (4)
where $H_{i}^{j}$ is the internal hidden state for the $j$-th token in layer $i$. In practice, we use Mamba (Gu & Dao, 2024) (a memory-efficient transformer alternative) followed by an MLP that projects features to the required dimension for our diffusion model. We find that a recurrence-based design elegantly handles long sequences of tokens, offering a strong balance between model capacity and memory usage. Later in experiments, we compare different recurrent architectures such as LSTM (Hochreiter et al., 1997) and causal-decoder transformers (Vaswani, 2017).
@@
Finally, we leverage a diffusion model to generate the actual parameter values. Drawing on p-diff (Wang et al., 2024) and MAR (Li et al., 2024a), we construct a 1D convolutional network that accepts random noise $\epsilon$ together with the prototypes $P$. Formally, at diffusion time step $t$, we aim to predict the noise $\epsilon$ in a denoising objective:
$L_{\text{diff}}\;=\;\mathbb{E}_{t,\,K,\,\epsilon}\bigl{\lVert}\,\epsilon\;-\;\epsilon_{\theta}\!\bigl{(}K_{t},\;t,\;P\bigr{)}\bigr{\rVert}^{2},$ (5)
where $\epsilon_{\theta}(\cdot)$ is our denoising network with learnable parameters $\theta$, and $K_{t}$ is the noisy version of $K$ at step $t$. Critically, the recurrent prototypes $P$ act as conditional inputs, enforcing global coherence among tokens. Gradients from this diffusion loss flow back into the recurrent model, implicitly learning token correlations as well.
Inference.
Once training is complete, the pipeline supports two inference modes. Basically, in conditional generation, we feed the permutation state $S$ of an existing checkpoint to replicate or perturb known parameter configurations. Further more, we can sample entirely unseen task states to produce novel parameter draws (Sec. 4). In both cases, the recurrent model outputs prototypes, and the diffusion model denoises random noise into a final parameter set that can scale up to hundreds of millions of weights.
Overall, recurrent diffusion elegantly decouples the complex problem of global parameter correlation into (i) a recurrent pass that focuses on cross-token interactions, and (ii) a diffusion pass that refines each token. This two-step design is precisely what allows RPG to handle large-scale architectures on standard hardware.
3 Main Experiments
@@ 3.1 Setup
We evaluate our method across various tasks, including ImageNet-1K (Deng et al., 2009) for theclassification, ADE20K (Zhou et al., 2017) for the semantic segmentation, COCO (Lin et al., 2014) for the object detection, and BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), and ARC (Clark et al., 2018) for the commonsense reasoning tasks. To verify the scalability, we conduct experiments on various architectures with parameter counts ranging from several to hundred million.
@@
We take parameters collection on the ImageNet-1K as an example. To save the cost, we finetune the full parameters of the models in timm  [^1] https://github.com/huggingface/pytorch-image-models and save 50 checkpoints as the training data. (More details are in Appendix C.6) For each checkpoint, we assign a unique permutation state to guide the generated parameters.
@@
Preprocessing and training details.
The length of parameter tokens, permutation states, position embeddings, and prototypes is set to 8192. It is worth noting that the permutation states and position embeddings are fixed during the training. We also study the influence of the token length, varying it from 1024 to 16384. The parameter diffusion consists of 1D convolutional layers. We default to using Mamba (Gu & Dao, 2024) as the architecture of the recurrent model. More details can be found in Appendix B.1.
@@
We input permutation states (relevant experimental results are in Appendix C.1) and position embeddings into the recurrent model to generate the prototypes. Then, the diffusion model utilizes the prototypes as conditions, along with random noises, to synthesize the entire network parameters. We repeat the above process 10 times and report the best, average, and minimum.